absolute value 28-2x=34
absolute value 9x-14=absolute value 7x +4
absolute value z/3-6=-6
absolute value 12n+4=0
absolute value 7+5c-14=-2

Answer by lynnlo(4176)

ANSWERS:
(1)X=-3
(2)X=9
(3)Z=0
(4)N=12
(5)C=1,-19/5